OpenStudy (anonymous):

What is the expression 12w9y3/-3w3y3 equivalent to

OpenStudy (anonymous):

\[\frac{ 12*w^9*y^3 }{ -3*w^3*y^3 }\]

OpenStudy (anonymous):

\[\frac{ 12 }{ -3 } * w^9*w^{-3} *y^3*y^{-3}\]

OpenStudy (anonymous):

-4*w^(9-3)*y^(3-3)=-4*w^6